### v1

# Microscopic Analysis on LLM players via Social Deduction Game

Abstract
Recent studies have begun developing autonomous game players for social deduction games using large language models (LLMs). When building LLM players, fine-grained evaluations are crucial for addressing weaknesses in game-playing abilities. However, existing studies have often overlooked such assessments. Specifically, we point out two issues with the evaluation methods employed. First, game-playing abilities have typically been assessed through game-level outcomes rather than specific event-level skills; Second, error analyses have lacked structured methodologies. To address these issues, we propose an approach utilizing a variant of the SpyFall game, named SpyGame. We conducted an experiment with four LLMs, analyzing their gameplay behavior in SpyGame both quantitatively and qualitatively. For the quantitative analysis, we introduced eight metrics to resolve the first issue, revealing that these metrics are more effective than existing ones for evaluating the two critical skills: intent identification and camouflage. In the qualitative analysis, we performed thematic analysis to resolve the second issue. This analysis identifies four major categories that affect gameplay of LLMs. Additionally, we demonstrate how these categories complement and support the findings from the quantitative analysis.
Index Terms:
Large language models, Natural language processing, Intelligent agents, Role-playing Games, Autonomous game players, Social deduction games
I Introduction
Conversation plays a pivotal role in social deduction games (SDGs), such as Among Us and Goose Goose Duck. In these games, players are divided into multiple adversarial groups and interact by executing actions through natural language sentences. Unlike other games, these interactions create tricky conversations where explicit statements often contradict the implicit strategies of the corresponding speaker. For example, in Among Us, imposters explicitly and repeatedly claim to be members of the crew. However, this assertion is part of their strategy to conceal their true identity. Consequently, crew members must uncover these hidden intentions to win. As this example illustrates, SDG players must possess the ability to navigate and manage such tricky communications.
Although such abilities are essential, traditional game agents may struggle to handle the tricky conversations occurring in SDGs. Recent studies have adopted large language models (LLMs) because LLMs are capable of generating fluent dialogues and strategic plans. Researchers believe that these capabilities could make LLMs suitable for supporting SDGs. However, most studies have primarily focused on improving gameplay performance rather than investigating whether LLMs can genuinely support these intricate conversations. Specifically, existing studies have two key issues when evaluating SDG-playing agents: (1) Macroscopic quantitative evaluation and (2) Non-systematic qualitative analysis.
First, the issue of macroscopic evaluation refers to the tendency of prior studies to evaluate various game-playing abilities using broad game-level outcomes instead of focusing on specific event-level skills. Existing research often relies on simple summative metrics, such as winning rates, which fail to account for the events that occur during gameplay[1, 2]. While these studies acknowledge the importance of certain skills for success in SDGs, they may overlook the need to measure these skills separately with such simple summative metrics. To scrutinize a specific skill, we suggest using microscopic analysis augmented with several specialized metrics. By performing a multifaceted formative analysis leveraging these metrics, one can reveal the true conversational abilities of LLMs in SDG play. In this study, we propose eight microscopic metrics to evaluate two essential skills: (1) Intent identification, the ability to combine explicit clues to uncover hidden information, and (2) Camouflage, which is the ability to disguise oneself from being detected as an opponent.
Second, the non-systematic evaluation issue indicates that previous studies lacked structured methods for error analyses. While certain studies identified abnormal patterns through simple qualitative analysis, the generalizability of their results was limited by their analytical procedures. For instance, some studies employed LLMs to analyze data without incorporating human-annotation procedures, while others did not adequately discuss their qualitative analysis methods for deriving generalizable abnormal patterns. To enhance the generalizability of qualitative analysis, we employ a systematic approach known as thematic analysis[3]. Moreover, we aim to draw meaningful insights by correlating the resulting categories with quantitative findings.
We conduct two studies to address these issues. In Study 1 (Section IV), we introduce microscopic metrics and discuss why macroscopic metrics are insufficient for analyzing specific skills. In Study 2 (Section V), we conduct a thematic analysis to derive categories and link these findings to the results of Study 1. In both studies, we let four LLMs play SpyGame, a variant of SpyFall with simplified actions (Section III-A). Consequently, this study makes the following contributions:
  • We demonstrate effectiveness of using microscopic event-level metrics to analyze performance of LLMs in SDGs. Specifically, we show that eight microscopic metrics offer a more detailed assessment than conventional macroscopic metrics in evaluating gameplay skills in SpyGame.
  • We conduct a qualitative analysis to identify categories of abnormal reasoning patterns in LLMs. Through thematic analysis involving multiple LLMs, we discover four main categories and five subcategories, including exposure, role ambiguity, memory distortion, and dissociation.
  • Finally, we validate our quantitative findings by correlating them with the results of the qualitative analysis.
II Related Works
This section reviews the existing literature in two ways. First, we provide a brief overview of LLMs and the historical attempts at developing LLM-based gameplay agents. Then, we point out their two issues on evaluating their agents: macroscopic evaluation and non-systematic analysis.
II-A Existing LLM-based Game Players
Before discussing LLM-based agents in SDGs, we first explain LLMs. An LLM is a language model trained on enormous corpora, excelling in various downstream tasks without requiring task-specific fine-tuning. Researchers in artificial intelligence have reported that LLMs exhibit human-like skills across diverse tasks, including question answering, communication, and planning [4, 5, 6]. For instance, LLMs have been employed to create chatbots that engage in natural, fluent conversations with humans[7]. Therefore, LLMs have been widely adopted in tasks involving natural conversations, such as humor, empathy, and decision-making[8, 9].
Recently, researchers have employed LLMs to design conversational agents for games. Leveraging their natural conversations, studies have reported that LLMs can play simple conversational games at a human level. For example, Brookins and Debacker[10] demonstrated that GPT-3.5 could play the Dictator game[11] and the Prisoner’s Dilemma game[12], displaying preferences for fairness and cooperation that often exceeded those observed in human players.
After confirming the gameplay ability of LLMs in simple conversational games, researchers have turned their attention to a more complex family of games: SDGs. SDGs require a higher level of communication skills like camouflage, which are not essential for playing simpler games. Therefore, researchers have developed autonomous player agents for SDGs, such as Avalon and Werewolf. For instance, Shi et al.[13] designed an LLM-based agent to play Avalon. To build working agents, they incorporated a reflection mechanism[14], which complemented the memory of LLMs with a human-inspired memory mechanism. Similarly, ReCon[15] offers additional hints to assist LLM players in choosing the best action during decision-making. Xu et al.[16] developed an LLM-based Werewolf player using reinforcement learning. They employed a success condition for specific actions as a reward signal to guide the learning procedure. However, these approaches have two key issues when evaluating the performance of their agents, which are discussed in the following section.
II-B Issues in Evaluating LLM Players
A fine-grained evaluation can point out weaknesses of an agent in the gameplay, which is crucial for designing effective LLM players. However, prior studies have typically employed more coarse-grained approaches. Therefore, these studies encounter two main issues in their evaluations: macroscopic quantitative evaluation and non-systematic qualitative analysis.
First, previous studies have leveraged macroscopic metrics for quantitative evaluation. Some researchers have adopted quantitative metrics, whose calculations are solely based on the results of entire game. For example, Xu et al.[2] evaluated LLM-based Werewolf players utilizing the winning rate, total game duration, and total number of camouflage behaviors. Similarly, Light et al.[17] employed winning rates and action success rates to assess gameplay, despite defining several specific gameplay skills, including deductive reasoning, coordination, collaboration, and deception. While these studies successfully measured the overall gameplay ability of an agent, their discussions remained superficial because the evaluation results did not support the existence of specific skills. Specifically, the winning rate (WR) does not provide insights into skills like deception or reasoning skills, as these skills all contribute to victory. Therefore, microscopic metrics are necessary to evaluate specific skills and conduct a multifaceted formative assessment of the gameplay of an LLM agent.
Second, prior studies have employed non-systematic qualitative analyses. Since macroscopic quantitative analysis alone cannot provide sufficient evidence for specific skills, some researchers have adopted qualitative error analysis to assess whether their agents possess certain skills[1, 17, 18]. For instance, Qiao et al.[1] provided a few samples for error analysis without presenting information on frequency of errors. Similarly, Kim and Kim[18] qualitatively analyzed errors made by LLMs in SpyFall. However, these errors are more related to generation error (e.g., contextual or format errors) than to specific gameplay skills. Though these approaches may provide insights into erroneous behaviors in LLM players, the generalizability of the results is questionable since the error may represent a specific but infrequent case. Also, other researchers have adopted language models to conduct quantitative analyses[19, 15]. For instance, Lan et al.[19] used ChatGPT[20] to assess agent gameplay across five aspects, including leadership and persuasion. While this approach can reduce the need for human labor in qualitative analyses, the reliability of these results is questionable due to potential issues with LLMs have reliability issues (e.g., hallucination) in generating answers. Hence, to obtain a comprehensive understanding of the errors made by LLM players, a systematic qualitative analysis of LLM gameplay is necessary.
In summary, the existing literature has two main issues. To address the first issue of macroscopic quantitative measures, we need microscopic measures that can investigate specific skills rather than evaluating overall gameplay ability. To tackle the second issue of non-systematic qualitative analysis, we need to adopt a systematic method for analyzing gameplay logs. In the following sections, especially in Studies 1 and 2, we demonstrate a method for conducting such required analyses using the SDG named SpyGame.
In summary, existing literature has two issues. To address the first issue of macroscopic quantitative measures, we need microscopic measures which can investigate a specific skills instead of whole gameplay ability. Also, to address the second issue of non-systematic qualitative analysis, we need to adopt a systematic way for analyzing gameplay logs. In the following sections, especially in Study 1 and Study 2, we will demonstrate a way to conduct such required analyses, using a SDG called SpyGame.
III Experiment
This section details the process of collecting gameplay logs from LLMs for both quantitative and qualitative analyses. To evaluate performance of vanilla LLM in SDG play, we let LLMs play an SDG through natural language interactions. Section III-A provides an overview of SpyGame, which is based on SpyFall. To collect the gameplay logs of SpyGame, we employed four LLMs, as detailed in Section III-B. The procedure for collecting logs is outlined in Section III-C.
III-A SpyGame
SpyGame is a modified version of SpyFall played by seven players. We altered Spyfall for two key reasons: First, LLMs may have been exposed to the rules and details of Spyfall during their pre-training phase. This exposure can cause LLMs to embed game-related knowledge from pretraining data in their parameters. Therefore, it is likely that these LLMs use this parametric knowledge as a gameplay hint rather than leveraging their inherent abilities. To mitigate this, we renamed the game to SpyGame.
Second, setting a time budget may influence the reproducibility of our analysis. Various factors in the experimental setup (e.g., the hardware used, network status, and model size) can impact inference time. Thus, to enhance reproducibility, we replaced the total time budget of a game with the total number of “turns” within a game.
III-A1 Team formation
Similar to SpyFall, SpyGame features two teams: six citizen players and one spy player. The citizens share a common location, called a hidden location, with each citizen plays a specific role in that location. However, the spy is not given information about the hidden location or the roles of the citizens. In addition, the citizens are unaware of the identity of the spy.
Players must uncover hidden information to win the game. In each turn, one player asks another player a question about the hidden location, and the questioned player must answer. Through these conversations, citizens try to identify the spy. Also, citizens should need to imply their knowledge of the hidden location without revealing too much detail, to avoid being suspected as the spy. Meanwhile, the spy attempts to determine the hidden location by pretending to be a citizen, gathering clues from the utterances of others. The detailed game rules, location-role sets, and prompts used for LLMs are provided in [anonymized for review].
III-A2 Game Flow
Each turn has a designated leader. The turn leader starts the turn by choosing one of two actions: questioning or accusation. If the leader selects the “question” action, they ask another player a question related to the hidden location. The questioned player must then respond based on their assigned role. Alternatively, if the leader opts for the “accusation” action, they pick the player whom the leader most suspects to be the spy and initiate a “Day Vote.” The other players, excluding the accuser and the accused, then vote on whether they agree with the accusation. If all players agree, the game ends immediately. After completing either the “question” or “accusation” action, the questioned or accused player becomes the leader for the next turn.
In every turn after the first, the spy player can publicly reveal their identity and guess the hidden location. Before each turn begins, the system asks the spy player if they are ready to reveal their identity. The game ends immediately when the spy player decides to reveal their identity. Note that in our implementation, instead of directly asking for the decision of the spy, we ask LLMs how certain they are about their guess.
Despite an accusation or an action of the spy, the game may continue until the final turn: Turn 9 in our implementation. At this point, the system initiates the “Final Vote.” All players secretly vote for the player they most suspect of being the spy. Following the vote, the game ends.
At the end of a game, the spy wins if any of the following conditions is met; otherwise, the citizens win:
  • A citizen is eliminated by a Day Vote.
  • The spy publicly reveals their identity and correctly guesses the hidden location.
  • The Final Vote results in multiple spy candidates, with more than one player receiving the most votes.
  • The Final Vote identifies only one spy candidate who is actually a citizen.
III-B Tested Models
We selected four LLMs for this study: GPT-4[4], GPT-3.5-turbo (Chat-GPT), Gemini-pro (1.0)[5], and LlaMA2-70b-chat[6]. We considered three criteria: model size, reasoning performance, and the existence of publicly-available API. Specifically, we selected models with a parameter size over 70 billion and a reasoning performance exceeding 70% accuracy[21] to ensure reasoning ability required for SDG play. Previous studies have reported a general performance trend among these four models: GPT-4 leads in performance, followed by Gemini, GPT-3.5, and LlaMA2.
We employed the following input and output procedures to analyze the vanilla performance of the four LLMs: First, as an input prompt, we provided only the minimal information necessary to play SpyGame. Specifically, the LLMs received inputs, including game rules, role descriptions, conversation history, and instructions for generating the next action. Unlike previous studies, we did not provide any additional information (such as hints) that could affect gameplay. Second, for output generation, we used a two-stage process involving free-form generation and JSON formatting. We requested LLMs to decide the next action in free form with their strategy. Then we instructed GPT-4 to rewrite the free-form output into a structured JSON format, which is compatible with our environment. By adopting these procedures, we attempted to minimize unintended effects stemming from additional information.
III-C Data Collection
To assess the gameplay of the spy players, we categorized citizens into two types: strong and weak. Since the game involves two adversarial teams, the performance of the spy can be influenced by the gameplay of the opposing citizens. If the citizen players are strong enough to quickly identify the spy, the spy may not have enough turns to demonstrate their abilities. Thus, we conducted separate experiments with weak citizens versus a spy and strong citizens versus a spy. Here, we used LLaMA2-70B-chat as the weak citizens and GPT-4 as the strong citizens, based on results from other reasoning tasks[21, 22, 4] and our preliminary experiments.
A total of 168 gameplay logs were collected. In this experiment, we adopted seven hidden locations from the original game. For each location, we ran 24 games: 2 types of citizens $\times$ 4 spy LLMs $\times$ 3 trials. Table I lists the statistics of the collected gameplay logs. On average, each game took 5.4 turns, with 23.8 reasoning steps. Here, the reasoning step means a unit of LLM’s reasoning for taking an action.
TABLE I: Average statistics for all players in the collected logs
Turns Reasoning Steps Utterances Words/Utterance
Spy vs. Strong 6.37 27.98 15.65 21.55
Spy vs. Weak 4.52 19.74 10.96 25.24
Total 5.45 23.86 13.31 23.07
IV Multifaceted Quantitative Analysis
Our study aimed to conduct a multifaceted quantitative analysis of LLM performance in SDGs. We designed seven metrics that enable fine-grained analysis of two key gameplay facets: intent identification and camouflage. Compared with previous studies that utilized single-faceted metrics, our approach allows for an analysis of intermediate gameplay behaviors. In this section, we discuss whether existing single-faceted metrics are sufficient for measuring intermediate behaviors. Next, we demonstrate a multifaceted quantitative analysis using the proposed metrics to explore gameplay behavior.
IV-A Insufficiency of Previous Analysis
Previous studies commonly adopted single-faceted analyses to examine SDG gameplay behavior[17, 16, 2, 1]. The most frequently used metrics are the winning rates (WR) and living rounds (LR). These metrics focus on the final state of the game to measure the performance of a player: WR utilizes winner information, while LR uses the number of turns played.
  • Winning Rate (WR) is the percentage of games won by the spy out of the total number of games.
  • Living Round (LR) is the average number of turns played across all games.
IV-A1 Metric result
Tables II (a) and III present the results of the single-faceted metrics when playing against strong and weak citizens, respectively. The WR and LR exhibited different trends in SDG play performance. In terms of WR, GPT-4 showed the highest performance against strong citizens (80.95%), followed by Gemini (57.14%), LLaMA2 (52.38%), and GPT-3.5 (33.33%). A similar trend was observed for weak citizens. However, in terms of LR, LLaMA2 demonstrated the highest performance against strong citizens (9.00), followed by Gemini (6.95), GPT-4 (5.80), and GPT-3.5 (3.71). This trend is similar when examining the performance against weak citizens. Thus, these metrics yield different results regarding the best-performing model: WR indicates GPT-4 as the best, while LR identifies LLaMA2 as the best.
This discrepancy is due to multiple winning conditions specified in the SDG. These conditions allow players to win games in several ways. Therefore, to accurately identify the skills that contribute to winning, it is necessary to consider multiple abilities simultaneously. However, since WR and LR consider only the final state of the game, these metrics cannot examine different abilities separately. For example, in SpyGame, the LR metric indicates how well a spy survives during the game. However, a longer survival duration does not always guarantee victory, leading to potential discrepancies between LR and WR results. Hence, a multifaceted method is essential for a comprehensive analysis of gameplay performance.
TABLE II: Quantitative results for playing against strong citizens
(a) General (b) Intent identification (c) Camouflage
Winning Living Guess Information Notice Caught Vote Vote
Model Rate (%) Round (#) Success (%) Deduction (%) Rate (%) Rate (%) Rate (%) Entropy
Gemini 57.14 6.95 50.00 30.00 33.33 022.22 35.7 0.348
(14) (20) (9) (42)
GPT-3.5 33.33 3.71 27.78 28.57 28.57 007.14 25.0 0.730
(18) (21) (14) (12)
GPT-4 80.95 5.80 81.25 61.90 61.90 025.00 33.3 0.953
(16) (21) (4) (30)
LlaMA2 52.38 9.00 - 00.00 09.52 100.00 45.2 0.634
(0) (19) (10) (126)
Total 55.95 6.365 35.78 30.86 33.33 037.83 40.47 0.642
Metrics whose lower values indicate better results.
TABLE III: Quantitative results for playing against weak citizens
(a) General (b) Intent identification (c) Camouflage
Winning Living Guess Info. Notice Caught Vote Vote
Model Rate (%) Round (#) Success (%) Catching (%) Rate (%) Rate (%) Rate (%) Entropy
Gemini 90.48 5.28 85.71 61.11 090.48 00.00 07.1 0.420
(14) (18) (2) (42)
GPT-3.5 71.43 2.52 71.43 92.86 076.19 $0$
(21) (14) (6) (0)
GPT-4 90.48 2.42 90.48 87.50 100.00 $0$
(21) (16) (2) (0)
LlaMA2 76.19 7.85 60.00 20.00 071.43 60.00 13.5 0.605
(5) (15) (5) (96)
Total 82.14 4.51 80.32 65.07 084.52 20.00 11.59 0.548
Metrics whose lower values indicate better results.
IV-B Intent identification
We designed multiple metrics to enable a multifaceted analysis of intent identification and camouflage. We operationally define Intent identification as the ability to uncover hidden information through deduction or by capturing pieces of information from the conversation. Intent identification has two sub-abilities: “information capture” and “information deduction.” Information capture mainly focuses on how well a spy grabs crucial hints that a citizen might inadvertently reveal. Since strong citizens rarely make such mistakes, it is highly challenging to evaluate this sub-ability against them. On the other hand, information deduction (ID) mainly focuses on how effectively a spy infers a hidden location by combining various pieces of information. Since weak citizens usually provide information that directly exposes the hidden location, it is not meaningful to assess this sub-ability against weak citizens. Therefore, metrics should consider whether the spy is effectively utilizing the information provided by these citizens. To evaluate both information capture and deduction, we analyzed outcomes of the guessing by the spy. The following subsections detail the design of the four metrics and their results. The results indicated that GPT-4 was the most powerful spy player, followed by Gemini, GPT-3.5, and LlaMA2.
IV-B1 Metric design
Based on a spy’s guesses during a game, we suggest four metrics for assessing intent identification. Four metrics range from relatively summative to highly formative.
  • Guess Success (GS) measures how much the guessing of a spy contributed to their victory. This metric focuses only on games where the spy attempted to guess, rather than all played games. We calculate GS by dividing the number of games in which the spy made a successful guess by the number of games in which guessing was attempted. A higher GS indicates that the spy has successfully identified hidden locations, demonstrating strong intent identification skills.
  • Notice Rate (NR) indicates how well a spy noticed the hidden location across all games. It is calculated by computing the percentage of games in which the spy correctly detected the hidden location over the total number of games played. Unlike GS, which is measured only when the spy decides to reveal their identity publicly, NR can be computed across all games. This is because NR mainly considers whether the spy “notices” the game regardless of whether they decide to guess or not.
  • Information Catching (IC) measures how often the spy successfully leverages exposed information to achieve a victory. Unlike NR, which only assesses the ability of the spy to identify the hidden location, IC evaluates the effectiveness in employing such information to win. We specifically focus on games where weak citizens directly exposed to information about the hidden location directly, instead of considering all played games. We measured IC by computing the proportion of games in which the spy successfully guessed the hidden location guessing from those where exposure occurred. Thus, a high IC score reflects the ability of the spy to achieve victory by effectively capturing and utilizing crucial information.
  • Information Deduction (ID) is similar to IC but it applies to games where exposure did not occur. This metric measures how well the spy combined pieces of information to determine the hidden location. To compute ID, we divided the number of games in which the spy successfully guessed the hidden location by the number of games without any exposure. Therefore, a high ID indicates that the spy successfully combined clues from the conversation and utilized them to achieve victory.
In particular, we devised a multifaceted interpretation of intent identification by using both IC and ID metrics simultaneously. By combining the IC and ID with their average values, three combinations were derived. The first combination is the most balanced, where both IC and ID are high. This represents an ideal scenario where the spy successfully identifies the hidden location regardless of the type of citizen. This signifies that the model is good at both sub-abilities: information capture and information deduction. The second combination occurs when only one of the metrics, either IC or ID, is higher than the average. This suggests a lack of proficiency in the corresponding sub-ability. For instance, a model with high IC but low ID demonstrates strong information capture but weak information deduction. The final and least favorable combination for intent identification is when both IC and ID are low. This combination indicates that the spy is unable to infer the hidden location in most games.
Moreover, by combining GS and NR, we can explore whether the spy effectively incorporates the incoming information to maintain their guess. The ideal combination is when both GS and NR are high. This indicates that the spy consistently maintained a correct guess after identifying the hidden location, until the end of the game. In contrast, a high GS paired with a low NR suggests that while the spy struggled to predict the hidden location initially, he/she manages to maintain their guess once the correct location is identified. Meanwhile, the combination of low GS and high NR implies that although the spy successfully guessed the hidden location, the spy is prone to discarding the guess easily. The worst-case scenario is when both NR and GS are low, indicating a significant failure in both identifying and maintaining the correct hidden location.
IV-B2 Result
Tables II (b) and III (b) present the measurements for intent identification when playing against strong and weak citizens, respectively. The results indicate that GPT-4 has sufficient intent identification ability compared to the other models. GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and LlaMA2 (81.25, 50.0, 27.78, and 0%, respectively). The results in terms of NR show a similar trend: GPT-4 (61.90%), Gemini (33.33%), GPT-3.5 (28.57%), and LlaMA2 (9.52%). Finally, regarding ID, GPT-4 was the best model, followed by Gemini, GPT-3.5, and LlaMA2 (61.90, 30.0, 28.57, and 0%, respectively). In summary, GPT-4 demonstrated the best performance against strong citizens in intent identification. Meanwhile, LlaMA2 showed the lowest performance against strong citizens.
Against weak citizens, GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and LlaMA2 (90.48, 85.71, 71.43, and 60%, respectively). The NR results displayed the same trend: GPT-4 (100%), Gemini (90.48%), GPT-3.5 (76.19%), and LlaMA2 (71.43%). Lastly, regarding IC, GPT-3.5 (92.86%) showed the best performance, differing in GS and NR, followed by GPT-4 (87.5%), Gemini (61.11%), and LlaMA2 (20.0%). Similar to the results against strong citizens, GPT-4 showcased good performance against weak citizens.
IV-B3 Discussion
In this section, we discuss how to derive insights using the proposed metrics to conduct a multifaceted intent identification analysis. This discussion encompasses three aspects: (1) a complementary relationship between IC and ID; (2) an exploration of the spy’s filtering behavior on incoming information; and (3) the need for formative metrics concerning intent identification.
First, ID and IC are complementary metrics that enable multifaceted analysis of intent identification. As aforementioned, these two metrics can be combined to form three distinct partitions, each representing different levels of intent identification ability. Using these partitions, we can determine the best model possessing sufficient sub-abilities for intent identification; In our analysis, GPT-4 demonstrated high scores in both ID and IC, with its ID being double that of the second-place model. Meanwhile, GPT-3.5 falls within the high-IC region but not the high-ID region. This indicates that while GPT-3.5 struggles to combine hints from the conversation to deduce the hidden location.
Second, by analyzing GS and NR, we can examine how the spy filters incoming information to maintain their guess. As suggested in the metric design, we can use four partitions made by GS and NR. Among the four models, GPT-4 belongs to the first partition and achieves a high score on both metrics. This indicates that GPT-4 can filter useful information, reinforcing the belief of the spy in the hidden location. This could be a hint that GPT-4 has strong information deduction sub-ability. Meanwhile, belonging to different partitions can induce differences in such filtering behaviors. For example, although Gemini and GPT-3.5 exhibited similar NR scores against strong citizens, the GS of Gemini was almost double that of GPT-3.5. This difference suggests that Gemini is more selective in using guess actions than GPT-3.5, as evidenced by the fact that Gemini used 14 guess actions regardless of citizens, whereas GPT-3.5 utilized 21 (against weak) and 18 (against strong) guess actions.
Finally, our intent identification metrics enable more detailed analyses of the factors contributing to the scores of simple metrics. Let us consider the simple metric, Win- ning Rate (WR), on Gemini models. Against strong citizens, Gemini showed a similar WR to LlaMA2. Meanwhile, against weak citizens, Gemini exhibited the same WR as GPT-4. However, these observations are somewhat shallow when we consider our intent identification metrics, as the three models display different intent identification abilities. For example, against strong citizens, Gemini had a higher ID score than LLaMA2 in that LLaMA2 completely failed to combine clues about the hidden location. Similarly, against weak citizens, we can conclude that GPT-4 is slightly better than Gemini because of its higher GS and NR scores. Thus, proposed metrics can reveal finer differences between LLMs than existing metrics.
In addition, such a detailed analysis revealed that GPT-3.5 is overly confident in its deduction results. As illustrated in Tables II and III, the spy with GPT-3.5 demonstrated relatively lower NR scores; however, the spy used guessing actions more frequently. That is, the spy rushed into making guesses even when their inference about the hidden location was incorrect. One possible explanation is that GPT-3.5 may treat its own inference as hard evidence, as discussed in Section V. We suspect that this mechanism may have contributed to GPT-3.5 achieving the lowest WR score, despite its relatively good intent identification ability.
IV-C Camouflage
In addition to intent identification, we designed four metrics for the multifaceted analysis of camouflage. Here, we operationally define camouflage as the ability to blend in with others during gameplay without revealing one’s true identity. Camouflage encompasses two sub-abilities: “evading suspicion” and “dispersing suspicion.” Evading suspicion mainly considers how well the spy avoids being suspected by others. It is quantified by the number of suspicions directed at the spy. On the other hand, dispersing suspicion primarily considers how difficult the spy makes it for all citizens to converge on a single suspect. It is quantified by the probability distribution of suspicions). To estimate the number and distribution of suspicions, we utilized logs of Day/Final votes. The following subsections describe the design of the four metrics and their results. The results indicate that GPT-3.5 is the most powerful spy player, followed by GPT-4, Gemini, and LlaMA2.
IV-C1 Metric design
Based on day/final votes and accusations during a game, we suggest four metrics for measuring camouflage, ranging from relatively summative to highly formative.
  • Caught Rate (CR) is how frequently the spy is caught by citizens in the Day/Final vote. We compute this by dividing the number of games in which the spy was caught by the number of games lost. Thus, a high CR indicates that the majority of citizens easily identified the spy player during the game. Note that this is a summative metric for the spy’s camouflage performance because it only considers the final state of the game.
  • Accusation Rate (AR) measures how frequently citizens accuse the spy. We measure this by computing the proportion of correct accusations made across all games. Unlike CR, which summarizes all citizens’ thoughts, AR focuses on whether an individual citizen can strongly suspect a true spy since the turn leader is the only player who can make an accusation. Depending on the gameplay, it is possible that there is no accusation during the game; in this case, AR cannot be measured.
  • Vote Rate (VR) is the proportion of suspicions about the spy across all votes that occurred. We measure this by computing the percentage of votes the spy received from citizens across the total number of votes, excluding the spy’s votes. Therefore, a high VR indicates that the spy failed to evade suspicion during the game.
  • Vote Entropy (VE) is how scattered the suspicions about the spy are in the game. We measure this by computing the entropy of the votes that citizens casted for each game and averaging them. Note that in SpyGame, a player’s suspicion cannot affect another player’s suspicion because they cannot have a discussion together. Thus, when a spy successfully blends with citizens, citizens have difficulty narrowing down the number of spy candidates. Thus, the value of VE increases.
We devised a multifaceted interpretation of camouflage by using VR and VE simultaneously. By partitioning VR and VE by their average values, we derived four combinations. First, the most balanced combination was a low VR and high VE. This is an ideal combination because the spy can successfully evade others’ suspicions and lead citizens to discourage each other from narrowing down spy candidates. Second, models in which VR and VE are both low or high lack one of the two sub-abilities. Finally, the worst camouflage combination occurs when VR is high and VE is low. This combination indicates that citizens’ suspicions were concentrated on the spy and that the spy failed to undermine suspicions.
IV-C2 Result
Tables II (c) and III (c) elaborate measurements of camouflage when playing against strong and weak citizens, respectively. We excluded AR from the tables because there was no accusation during gameplay, regardless of the citizen type. Similarly, some values are missing for VR and VE because the spy ended the game early without votes by guessing the hidden location.
The results indicate that the variants of GPT have sufficient camouflage ability compared to other models. Against strong citizens, GPT-3.5 was the best model in terms of CR, followed by Gemini, GPT-4, and LlaMA2 (7.1, 22.2, 25.0, and 100%, respectively). The results for VR showed a similar tendency: GPT-3.5 (25%), GPT-4 (33.3%), Gemini (35.7%), and LlaMA2 (45.2%). Finally, regarding VE, GPT-4 was the best model, followed by GPT-3.5, LlaMA2, and Gemini (0.953, 0.730, 0.634, and 0.348, respectively). Interestingly, GPT-3.5 showed camouflage performance similar to that of GPT-4, although GPT-4 showed much higher performance in other language-based tasks[4].
Against weak citizens, the GPT variants showed good performance compared to others. Because the GPT variants ended the game by guessing, citizens had no chance to find the spy. Thus, the CR values for GPT-3.5 and GPT-4 were zero, and VR and VE could not be measured using these models. Other models showed a similar tendency to that against strong citizens. Regarding CR and VR, Gemini (0% and 7.1%, respectively) was better than LLaMA2 (60% and 13.5%, respectively). However, regarding VE, Gemini (0.420) performed worse than LLaMA2 (0.605), as shown by the results against strong citizens.
IV-C3 Discussion
In this section, we discuss how to derive insights using the proposed metrics to conduct a multifaceted camouflage analysis. This discussion includes the following two aspects: (1) a complementary relationship between Vote Rate (VR) and Vote Entropy (VE) and (2) the need for formative camouflage metrics.
First, VR and VE are complementary metrics that enable a multifaceted analysis of camouflage. As suggested above, these two metrics can form three partitions explaining the different levels of camouflage ability. Using these partitions, we can determine the best model for balancing the two sub-abilities of camouflage; the GPT family showed gameplay balanced between VR and VE. These models can evade and disperse suspicion to win the game. In contrast, the worst model in terms of camouflage was LLaMA2. Although LLaMA2 performed better against weak citizens than strong citizens, LLaMA2 player was often suspected as a spy and disregarded the fact that the majority of players doubted him/her. Here, the two sub-abilities are both good or both bad in these two cases. Gemini showed a different behavior: the model was good at evading suspicion but struggled to disperse suspicion. Thus, as illustrated above, VR and VE can provide a wider picture of camouflage behavior.
Second, the proposed metrics provide additional perspectives for investigating camouflage compared to simple summative metrics. In Section IV-A, we observed a discrepancy between two summative metrics, winning rate (WR) and living round (LR). For example, both GPT-3.5 and LLaMA2 showed relatively low WR, but the performance gap regarding LR between them is large. Therefore, using only summative metrics, we can conclude that LLaMA2 outperformed GPT-3.5 in SpyGame. However, our metrics suggested different results; GPT-3.5 is good at camouflage, but LLaMA2 is not. This example shows that the proposed metrics are more suitable for examining camouflage ability than existing summative metrics.
Second, proposed metrics provide additional perspectives on investigating Camouflage, compared to simple summative metrics. In section IV-A, we observed a discrepancy between two summative metrics, winning rate (WR) and living round (LR). For example, both GPT-3.5 and LLaMA2 showed relatively low WR scores, but regarding LR, achieved relatively low and high performance, respectively. So, using summative metrics only, we may conclude that LLaMA2 outperforms GPT-3.5 in SpyGame. However, the proposed metrics suggested different results; GPT-3.5 is good at Camouflage, but LLaMA2 is not. This example shows that proposed metrics are more suitable to examine Camouflage ability than existing summative metrics.
Similarly, camouflage metrics are helpful in identifying gameplay problems. For example, both GPT-4 and GPT-3.5 showed relatively low LR scores; however, WR achieved the best and worst scores, respectively. Therefore, using only summative metrics, we can conclude that GPT-4 outperforms GPT-3.5 in SpyGame. However, the proposed metrics suggest different results: both GPT-4 and GPT-3.5 are good at camouflage regarding VR and VE.
V Qualitative Analysis
To reconfirm the findings of the quantitative analysis from another perspective, we conducted a thematic analysis. We defined three types of reasoning errors in LLM reasoning while deciding their actions by coding abnormal patterns of 15 randomly selected gameplay logs. As a result, we found that these patterns are related to psychological concepts, including identity confusion, memory distortion, and dissociation[23].
In addition, we found that these concepts are related to the results of the quantitative analysis, especially for caught rate (CR) and guess success (GS). The following subsections illustrate the procedure and results of this qualitative analysis.
V-A Analysis Procedure
Through thematic analysis, we labeled the abnormal patterns of players in the gameplay logs. To enrich our quantitative analysis results, we adopted a bottom-up inductive approach to discover the underlying patterns in the data. The detailed analysis process was as follows. First, we defined the initial labels of the abnormal patterns from 15 randomly selected gameplay logs. Three annotators with computer science backgrounds participated in this process. They read the logs to understand the spy’s gameplay situation when establishing initial labels. To avoid bias towards a specific LLM, we hid the information of the models in the process. Second, we refined these initial labels through an iterative process. Because LLMs perform one or two reasoning steps to decide on an action during the gameplay of SpyGame, a gameplay log is a series of several reasoning steps. So, we used one step as the unit for this analysis. We repeated this process five times between December 29, 2023 and June 1, 2024 until agreement on the identified labels was saturated. The agreement was measured using Fleiss’ Kappa[24], and the final agreement between the annotators ranged from 0.486 to 0.646, which indicates moderate agreement. Finally, using these settled labels, we labeled 506 reasoning steps from 168 games. Consequently, we identified three major categories and five subcategories.
TABLE IV: Frequency of each category found in qualitative analysis
$N$ Role ambiguity Memory distortion Dissociation
Against weak citizens
Gemini 61 1.63 4.91 3.27
GPT-3.5 35 2.85 2.85 8.57
GPT-4 34 0.00 0.00 5.88
LlaMA2 77 9.09 6.49 9.90
Against strong citizens
Gemini 82 3.65 6.09 15.68
GPT-3.5 39 0.00 10.25 23.07
GPT-4 71 4.22 4.22 9.85
LlaMA2 107 9.09 10.28 14.95
The total number of reasoning steps in each condition
V-B Resulting Categories
This section illustrates the results of our thematic analysis. For each category, we explain its definition, report its frequency, and provide some examples. Table IV summarizes the frequencies of all categories.
V-B1 Exposure
Exposure is a case in which a citizen’s utterance literally reveals the hidden location. The Fleiss’ Kappa value for this category was 0.7875. Note that this pattern can only be identified in utterances rather than in reasoning steps. Therefore, we report the frequency of this category based on utterances; among 729 citizen utterances, 125 utterances were labeled as exposure. The following quote provides a representative example:
[The hidden location is an airplane]
LlaMA2 (Citizen, 1st-class passenger):
‘‘What is the maximum speed of an airplane in an emergency situation?’’
We suspect that this exposure pattern is because LLMs may not have sufficient common sense to understand the game rules. In particular, SpyGame has a rule “citizens should not expose the location to the spy.” A human citizen player can notice that a direct mention to other players of the hidden location during a game causes their defeat. However, some LLMs may not have this common sense among players; they simply go through the mentions. Therefore, LLMs believe that they can reveal their information to other citizens because the game rule does not prohibit such behaviors and the citizens already know that fact. Thus, citizen LLMs sometimes publicly reveal their hidden information.
V-B2 Role ambiguity
Role ambiguity[25] is a case in which the spy confuses their own identity with that of a citizen. In an SDG, understanding a player’s identity is related to two major concepts affecting gameplay: (1) the team of the player and (2) the goal of the player. Therefore, we divided abnormal patterns of role ambiguity into two sub-categories: team misunderstanding and goal misunderstanding. These two subcategories can occur independently. We present details of the subcategories in the below paragraphs. In summary, Fleiss’ Kappa for this category was 0.563. Among 506 reasoning steps in the data, we identified 21 steps with role ambiguity.
Team misunderstanding
This is a case in which the spy disbelieves that they belong to citizens. Although this can occur in any reasoning step, we provide one example here:
[The spy tries to answer a citizen’s question. Before that, the spy generates its reasoning first.]
Gemini (Spy): Emphasize the importance of ambiance and personalized service, while avoiding giving away any specific details about the location...
Goal misunderstanding
This is a case in which the spy misbelieves that their goal is the same as that of the citizens. Here is an example:
[The spy tries to decide the next action. After selecting ‘question’ as the next action, the spy generates its reasoning.]
LlaMA2 (Spy): ... By asking this question, I hope to gain more information about the location and potentially identify the spy. If the spy is unaware of the restaurant’s cuisine, ...
We suspect that this category occurs because of the memory limitation of LLMs. To understand a player’s identity correctly, an LLM should retain information about the player’s identity during the decoding process. Since we provided the game rule, identity of a player, and conversation logs together as an input prompt, the LLM needs to distinguish information about the identity from the input prompt. However, such a distinction is difficult for the LLM because the other parts of the input prompt are much larger than the identity part. For example, the explanation of the game rules takes more than 130 words, whereas that of identity takes only 10 words. Thus, LLM players can confuse their identities with other characters written in the game rules, triggering role ambiguity.
V-B3 Memory distortion
Memory distortion occurs when a spy uses false or distorted memory to decide their next action. For this category, the Fleiss’ Kappa value was 0.486 among the three annotators. We identified 32 steps with memory distortion among 506 reasoning steps from the data. An example is as follows:
[Previously, Player2 and Player3 did not have any conversation. The spy tries to guess the hidden location.]
GPT-3.5 (Spy): Based on the conversation between Player2 and Player3, it seems that the location is likely ...
We suspect this category is identical to the common “hallucination” phenomenon of LLMs, where a language model generates non-sensical or unfaithful text for the provided source input[26]. In a similar vein, our memory distortion makes the LLM use inappropriate, false, or distorted clues to decide the next action or utterance. These actions, which are affected by hallucination, can hinder the gameplay.
V-B4 Dissociation
Dissociation is a case in which the spy employs their own utterances as evidence of the hidden location because they misunderstand the source of the utterance as other players. The annotators agreed on this category with a Fleiss’ Kappa value of 0.646. In this category, we identified 59 of 506 reasoning steps in the data. An example is as follows:
[Player4, the spy, tries to guess the hidden location.]
LlaMA2 (Spy): Based on the conversation, I would say that the location is most likely a workplace or office. Player4 described their environment as well-lit, well-ventilated, ...
We suspect that this category occurs because LLMs are not good at differentiating between numerical symbols. As previous studies have reported that LLMs suffer from biases induced by proper names[27], we named each player as ‘Player#’ (e.g., Player1) to avoid such biases. Therefore, when an LLM attempts to collect clues from the conversation log, it has to distinguish the source of each utterance using a single difference between player names: a single digit. However, previous studies have reported that LLMs struggle to recognize different digit tokens during their computation processes. Thus, LLM players can misinterpret their words as those of others in the conversation log, which triggers dissociation.
V-C Reconfirmation of Quantitative Analysis
Combining the results of the two analyses confirms the findings of the quantitative analysis. Specifically, we discovered that these quantitative labels, except exposure, can be connected to two abilities in the quantitative analysis: camouflage and intent identification.
First, role ambiguity increases the likelihood of a spy’s camouflage failure. When we compared the frequency of role ambiguity and the results of the three camouflage metrics, the trend of frequency was similar to that of the metrics. For example, against strong citizens, GPT-3.5 showed the lowest errors in role ambiguity and achieved the highest performance in camouflage metrics. Similarly, LLaMA2 showed more errors while achieving a lower performance than the other models. We suspect that this similarity between the trends is due to the loss of the spy’s need for camouflage. When role ambiguity occurs in the spy’s reasoning steps, the spy confuses their role with that of a citizen. Such confusion leads the spy to make an unusual utterance, which draws suspicion.
Second, memory distortion and dissociation may hinder a spy’s intent identification during the game. When we compared the frequency of memory distortion and dissociation with the results of the four intent identification metrics, the trend of frequency was similar to that of the metrics. For example, against strong citizens, GPT-3.5 showed the highest errors in both memory distortion and dissociation and achieved the lowest performance in intent identification. Similarly, GPT-4 showed fewer errors while achieving a higher performance than the other models. We suspect that this similarity between trends is due to the quality of information that the spy considers during their reasoning process. The spy player should select and combine pieces of information from the conversation to identify the hidden location. When memory distortion occurs, the spy can regard misinterpreted pieces as appropriate because they distorted the information. Similarly, when dissociation occurs, the spy regards the apocryphal information that they previously said as information provided by citizens. Therefore, such misunderstood pieces can be obstacles to reaching the correct answer.
VI Conclusion
To assess whether LLMs can effectively support tricky conversations used in SDGs, we tackled two key issues with previous evaluation methodologies. The first is the reliance on macroscopic quantitative evaluations, which overlook the need for microscopic formative assessment within quantitative analyses. The second is the lack of non-systematic qualitative evaluation, highlighting the necessity for thematic analysis to uncover the logical reasoning behind gameplay. In our demonstration, we chose SpyGame, a variant of SpyFall, as the experimental environment. While human gameplay often incorporates nonverbal and paraverbal cues to identify the hidden intention of others, we adopted the simplest implementation of LLM players to simplify our demonstration. We analyzed the gameplay behaviors using four different LLMs in SpyGame logs. Our analysis was conducted in two ways. First, through quantitative analysis, we employed eight metrics to address the macroscopic evaluation issue. This analysis revealed that our metrics are more effective than existing macroscopic metrics for explaining intent identification and camouflage behavior. Second, in qualitative analysis, we performed thematic analysis to resolve the non-systematic evaluation issue. From the analysis, we discovered four major categories that affect LLMs’ gameplay as a spy. These major categories supported the findings of our quantitative analysis. Despite the limitations of our study, our work underscores the significance of microscopic analysis in evaluating LLMs’ gameplay. In the future, we plan to extend this work to multi-modal LLMs to better capture the differences between human and LLM players. We hope that this study provides valuable insights into the behavior of LLM-based players in SDGs.
VII Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2021-0-01341, Artificial Intelligence Graduate School Program, Chung-Ang University) and the Chung-Ang University Graduate Research Scholarship in 2024.