### v2

Abstract
Recent studies have investigated whether large language models (LLMs) can support obscure communication that requires specialized skills, such as inferring subtext or doublespeak. To conduct the investigation, researchers have used social deduction games (SDGs) as their experimental environment, in which players conceal and infer specific information. However, prior work has often overlooked how LLMs should be evaluated in such settings. Specifically, we point out two issues with the evaluation methods they employed. First, metrics used in prior studies are coarse-grained as they are based on overall game outcomes that often fail to capture event-level behaviors; Second, error analyses have lacked structured methodologies capable of producing insights that meaningfully support evaluation outcomes. To address these issues, we propose a macroscopic and systematic approach to the investigation. Specifically, we introduce seven fine-grained metrics that resolve the first issue. To tackle the second issue, we conducted a thematic analysis and identified four major reasoning failures that undermine LLMs’ performance in obscured communication.
Index Terms:
Large language models, Natural language processing, Intelligent agents, Role-playing Games, Autonomous game players, Social deduction games
[…]
=-21pt
I Introduction
Large language models (LLMs) have shown that they can successfully conduct typical informative communication with humans by leveraging their natural language understanding abilities [1, 2, 3]. Inspired by these reports, there is a growing interest in adapting LLMs for obscured communication as a next step [4]. Here, the obscured communication indicates that the intended meanings of messages present in the communication are obscured or made unclear to understand, usually by controlling the spread of information. For example, in the real world, an imaginary role-play requires obscure communication to boost immersion in such a role-play. Among the required skills, doublespeak is one of the representative skills that enable a participant to achieve an advantageous position in such a play. So, since LLMs with doublespeak ability can support the play, researchers have recently focused on investigating such abilities. [5]
Meanwhile, investigating whether LLMs can conduct obscure communication requires a specialized environment. As ordinary LLM use cases focus on providing information, obscured communication cannot be tested with such ordinary cases. To set a more suitable environment for the examination, researchers have used social deduction games (SDGs) because SDGs force players to use doublespeak to achieve their goals. Specifically, they let LLMs play an SDG and collected the game logs, including LLMs’ decisions and conversations. After, they devised several metrics using particular game actions for evaluating LLMs’ ability of doublespeak. [6] Also, they conducted qualitative analyses to discover erroneous actions or utterances. For example, [7] examined whether LLMs can support obscured communication by letting them play a famous SDG, Werewolf. However, to thoroughly examine LLMs’ doublespeak abilities, these attempts have two issues: (1) Macroscopic quantitative evaluation and (2) Non-systematic qualitative analysis.
First, the issue of macroscopic evaluation refers to the tendency of prior works to use broad game-level outcomes as their metrics instead of specific event-level actions. Existing research often relies on simple metrics, such as win rates, which fail to account for the events that occur during gameplay [6, 7]. While these studies acknowledge the importance of certain skills for supporting obscured communication, they may overlook the need to measure these skills separately with such granular metrics. To scrutinize a specific skill, we suggest using microscopic analysis augmented with several fine-grained metrics. Specifically, we propose seven metrics to evaluate two essential skills regarding doublespeak: (1) Subtext inference, the deductive ability to identify doublespeak of others, and (2) Deceptive control, which is the ability to control information to make others confused.
Second, the non-systematic evaluation issue indicates that previous studies lacked structured methods for error analysis. While certain studies identified abnormal patterns through simple qualitative analysis, the reproducibility of their results was limited by their analytical procedures. For instance, some studies employed LLMs to analyze data without incorporating human-annotation procedures [8, 9], while others did not adequately discuss their qualitative analysis methods for defining abnormal patterns. Also, they simply reported the occurrence tendency of the patterns rather than discussing how they affected their evaluation results [10]. We believe that conducting qualitative analysis in a systematic way provides both reproducibility and interoperability of experimental results. Therefore, we adopt a systematic approach known as thematic analysis [11] to identify recurring reasoning errors that hinder LLMs’ doublespeak. By correlating these qualitative patterns with our quantitative metrics, we uncover how specific reasoning failures impact LLMs’ performance in subtext inference and deceptive control.
We conducted two studies to address these issues. In Study 1 (Section IV), we perform a quantitative analysis using our fine-grained metrics. The result reveals how coarse-grained metrics have overlooked finer aspects of doublespeak. In Study 2 (Section V), we conduct a thematic analysis on collected logs. Based on the result, we deduce several categories and link these findings to the results of Study 1. In both studies, we let four LLMs play SpyGame, a variant of simple SDG called SpyFall (Section III-A). Consequently, this study makes the following contributions:
  • We demonstrate the effectiveness of using microscopic quantitative analysis to assess the doublespeak of LLMs. Specifically, using seven fine-grained metrics, we show possibility of studying doublespeak in LLMs’ SDG play.
  • We identify four categories of abnormal reasoning patterns that harm obscured communication. Also, by combining results from the two studies, we discover a correlation between the results of the two analyses.
II Related Works
This section reviews the existing literature in two ways. First, we outline how recent research has shifted its focus from evaluating LLMs on informative communication tasks to more complex settings that require obscured communication. Then, we point out two issues posed in their assessment: macroscopic evaluation and non-systematic analysis.
II-A Adapting LLMs for Obscured Communication
LLMs have shown human-level performance in various informative communications. For example, researchers widely tested LLMs’ ability to question-answering or summarization [12, 13]. To accurately and reliably provide answers, LLMs for informative communication attempt to leverage true knowledge and suppress false statements during the generation process [14, 15]. So, researchers attempt to build an honest and harmless assistant for informative communication.
Noting that real-world communication includes both informative and obscured communication, researchers have recently turned their interests to obscured communication. The two types of communication differ in how each participant views their opponents. In informative communication, participants trust their opponents and try to establish a cooperative relationship. Meanwhile, participants lack trust in their opponents in obscured communication; thus, they seldom provide valuable information to others and often form adversarial relationships. So, participants require different abilities to achieve their goals in such an adversarial situation; skills like doublespeak are needed. Specifically, participants should manipulate how they present the information they have or capture subtext from the opponent’s utterances.
In developing negotiation or debate situations, the ability to discern the opponent’s intention in obscured communication is crucial, as subtext is often used in such communication. For example, in the context of applying LLMs to negotiation, prior research has investigated how to measure the ability of an LLM to infer the preferences of the opponent by identifying hidden intentions and meanings in the conversation [16]. Another example research applied LLMs to debates in order to identify and leverage weaknesses in the opponent’s arguments. Here, LLMs’ ability to discern intentions plays a crucial role, helping to accurately recognize hidden intentions or logical weaknesses in the opponent’s claims [17].
As obscured communication can be naturally seen in games, researchers have recently evaluated relevant skills using social deduction games (SDGs). As SDGs force players to utilize obscured communication to win the game, an SDG can be a suitable testbed for assessing or quantifying such communication. So, researchers aimed to demonstrate how well their agents can do obscured communication by letting them play SDGs. For example, Wang et al. [9] proposed an LLM agent capable of playing the Avalon game, introducing mechanisms such as dual contemplation to control conversational information. Similarly, Shi et al. [18] proposed another LLM agent that can play Avalon. Other than these zero-shot attempts, researchers used training methods to improve their agents. For instance, Xu et al. [19] trained an LLM agent applying reinforcement learning to play the Werewolf game. Collectively, these works demonstrate that SDGs offer a structured and adversarial context in which LLMs’ handling of obscured communication can be rigorously assessed.
II-B Attempts to assess LLMs in obscured communication
To enhance LLMs’ ability to support obscured communication, it is essential to identify the specific skill areas that require improvement. Therefore, most researchers have conducted evaluations in both quantitative and qualitative ways to reveal the current status of LLMs in obscured communications. For example, Xu et al. [7] and Light et al. [20] proposed automated methods for quantitatively evaluating their LLM-based players who play Werewolf or Avalon games. Specifically, they adopted particular game conditions, such as win rate or action success rate, that occur during their environmental games as their evaluation metrics.
However, we suspect that such metrics are too coarse-grained to adequately capture intermediate behaviors during gameplay. Specifically, the win rate is only available post hoc, limiting its utility for formative assessment. Additionally, the action success rate may fail to account for the multi-player dynamics of SDGs in which performance involves not only task execution but also managing others’ perceptions, such as avoiding suspicion. In other words, achieving a higher win rate or action success rate in an SDG does not entail that the model possesses superior competence across all sub-skills involved in obscured communication.
Other researchers proposed more granular metrics, such as logicality or social behaviors, that can be computed by scoring the reasoning or statements made by LLMs during an Avalon game[9, 8]. However, such metrics proposed in those studies rely on an auxiliary LLM to qualitatively assess reasonings or statements, rather than being grounded in objective in-game events. Therefore, fine-grained quantitative metrics grounded in objective in-game events are essential for evaluating LLMs’ abilities in multi-agent obscured communication, such as subtext inference and deceptive control.
While quantitative evaluation offers valuable insights into measurable performance, it is often insufficient to fully capture the underlying reasoning errors or abnormal behaviors exhibited by LLMs. To address this limitation, researchers have conducted qualitative analyses either as a complement to quantitative evaluation or as a standalone method. Kim [10] presented the frequency of misbehavior of LLMs while playing SpyFall to support their assessment of LLMs in the game. Quia et al. [6] presented some particular cases of LLMs to support their assessment results driven by their evaluation framework using communication games.
However, the methods used in their qualitative analyses were less structured or not related to quantitative results. For example, the categories adopted in [10] related to the generation procedure of LLMs, such as formatting errors, rather than LLMs’ ability to support their obscured environment. To enhance the utility of qualitative analysis, it should serve not only to validate the outcomes of quantitative evaluations but also to provide insights into the underlying causes of results.
In summary, the existing literature has limitations in its quantitative and qualitative analysis methods. To discover whether LLMs can support obscured communication, quantitative analysis with fine-grained metrics and validation of qualitative analysis are necessary.
III Experiment
This section details the process of collecting gameplay logs from LLMs for both quantitative and qualitative analyses. Similar to prior works, we collected gameplay logs of LLMs by letting play SpyGame, which is suitable for observing obscured communication. Section III-A provides an overview of SpyGame, which is based on Spyfall. We employed four LLMs to collect the gameplay logs of SpyGame. Section III-B details the tested models and procedure of data collection.
III-A SpyGame
Figure 1: SpyGame
SpyGame is a variation of Spyfall played by seven players [^1] https://github.com/elu-lab/spygame. To better assess LLMs’ capacity for obscured communication, we modified the original Spyfall in two ways. First, we renamed the game to SpyGame, rather than using Spyfall. LLMs may have been exposed to the rules and details of Spyfall during their pre-training phase. This exposure can cause LLMs to embed game-related knowledge, such as rules or hints, from pretraining data in their parameters. Therefore, it is likely that these LLMs exploit memorized game-specific knowledge from pretraining, rather than demonstrating reasoning-based communicative skills during gameplay. So, to avoid such pretrained knowledge, we changed the name of the game to SpyGame.
Second, we replaced the total time budget of a game with the total number of “turns” within a game. Setting a time budget may influence the reproducibility of our analysis. Various factors in the experimental setup (e.g., the hardware, network status, and model size) can impact inference time. Thus, to enhance reproducibility, we used the number of turns.
III-A1 Team formation
Similar to Spyfall, SpyGame requires two teams: six citizens and one spy. The citizens share a common location, called a hidden location, with each citizen playing a specific character in the location. For example, let us assume that the location is a school. Then, the citizens are assigned to distinct characters, such as a student or a principal. Meanwhile, the spy does not have any information about the location or characters.
Both teams must manage what information they reveal and conceal throughout the game. Citizens should protect the information about the location, and the spy should pretend to be a citizen. So, to win the game, players must uncover hidden information. Simultaneously, all players should avoid being suspected of being a spy. Each citizen’s utterance should reveal just enough to signal alignment with other citizens, while minimizing the risk of exposing the hidden location to the spy. Meanwhile, the spy should gather subtexts from the conversation and deduce the hidden location.
This indicates that the abilities required of both the citizen and the spy are centered around concealing and revealing specific information. Meanwhile, because SpyGame is adversarial in nature, a substantial imbalance in competence between the evaluated player and their opponents can compromise the validity of the evaluation. Moreover, when evaluating a citizen player, the influence of fellow citizens can confound the measurement, making it more difficult to isolate the individual player’s ability. For these reasons, we restrict our evaluation to the spy, who operates somewhat independently. To assess the performance of spy under varying team conditions, we constructed two scenarios: one with strong citizens and one with weak citizens.
III-A2 Game Flow
SpyGame progresses over multiple turns, with each turn consisting of a specific game event. Figure 1 illustrates the four key events: (a) questioning and answering, (b) accusation and day vote, (c) guessing, and (d) final vote. In each turn, a designated leader initiates the turn by choosing one of two actions: questioning or accusation. If the leader chooses to question, they ask another player about the hidden location (Figure 1a). Then, the questioned player responds based on their assigned character.
If the leader opts for an accusation, they point out the player they suspect to be the spy, triggering a Day Vote (Figure 1b). All players except the accuser and the accused participate in the vote. If all players agree, the game ends immediately. After completing either questioning or accusation, the questioned or accused player becomes the leader for the next turn.
From the second turn, the spy player guesses the hidden location by herself (Figure 1c). For each turn, she tells her guess to the system secretly and provides a number indicating how certain she is. When the certainty is high enough [^2] In our implementation, we use 9 as the threshold out of 10., the system announces her guess and ends the game. Otherwise, the game continues.
To avoid endless play, we set a final turn [^3] In our implementation, we set the last turn as 9.. When players finish the final turn, the “Final Vote” begins (Figure 1d). All players secretly vote for the most suspicious player in their minds. At the end of a game, the spy wins when one of the following conditions is satisfied. Otherwise, the citizens win.
  • All players agreed to eliminate a citizen in a Day Vote.
  • After the Final Vote, a citizen has the highest vote without a tie.
  • The spy’s guess is correct when the system announces the spy’s guess.
III-B Models and Data
We selected four LLMs for our experiment: GPT-4[1], GPT-3.5-turbo (ChatGPT), Gemini-pro 1.0[3], and Llama2-70b-chat[2]. As obscured communication requires reasoning on subtexts, we selected large enough models with sufficient reasoning ability on simple tasks. Specifically, we selected models with a parameter size over 70 billion and a reasoning performance exceeding 70% accuracy[21] in 2023. LLMs provide their actions or utterances with their reasoning in natural language. So, we employed GPT-4 to sanitize their outputs into a machine-readable format.
To collect data, we let each tested model play SpyGame as a spy against strong or weak citizens. Here, the strong or weak citizen models were fixed based on performance reported in prior studies [21, 22, 1]: GPT-4 for strong and Llama2-70b-chat for weak citizens. And, we conducted our experiment repeatedly. Specifically, we used seven locations borrowing from the original game. For each location, we ran 24 games4442 types of citizens $\times$ 4 spy LLMs $\times$ 3 trials.. Consequently, we collected a total of 168 gameplay logs. Table I lists statistics of collected logs. On average, each game took 5.4 turns, with 23.8 reasoning steps which refers to a unit of LLM’s reasoning for taking an action.
TABLE I: Average statistics for all players in the collected logs
Turns Reasoning Steps Utterances Words/Utterance
Spy vs. Strong 6.37 27.98 15.65 21.55
Spy vs. Weak 4.52 19.74 10.96 25.24
Total 5.45 23.86 13.31 23.07
IV Microscopic Quantitative Analysis
Our study aimed to investigate whether LLMs can support obscured communication by conducting a microscopic analysis of their SDG playing. We designed seven metrics that enable fine-grained analysis of two key SDG facets: subtext inference and deceptive control. Compared with most metrics adopted by previous research, which were designed on game-level outcomes, our approach allows for an analysis of intermediate behaviors during the game. Also, our quantitative metrics are fully automated metrics that do not require any manual labeling. In this section, we first discuss whether existing simple metrics are sufficient for measuring intermediate behaviors of LLMs. Then, we demonstrate our analysis can discover new insights that could not be captured by using current metrics by discussing our quantitative results.
IV-A Insufficiency of Previous Analysis
Previous studies commonly adopted macroscopic analyses to examine SDG gameplay behavior[8, 20, 19, 7, 6]. The most frequently used metrics are the win rates (WR) and living rounds (LR). These metrics evaluate a player’s performance based on final game outcomes: WR is determined by the win/loss result, while LR reflects how long the player remained active during the game.
  • Win Rate (WR) is the percentage of games won by the spy out of the total number of games.
  • Living Round (LR) is the average number of turns the spy survived, across all games.
IV-A1 Metric result
Tables II (a) and III (a) present the results of the coarse-grained metrics when playing against strong and weak citizens, respectively. The WR and LR exhibited different trends in SDG play performance. In terms of WR, GPT-4 showed the highest performance against strong citizens (80.95%), followed by Gemini (57.14%), Llama (52.38%), and GPT-3.5 (33.33%). A similar order of performance was observed for against weak citizens. However, in terms of LR, Llama demonstrated the highest performance against strong citizens (9.00), followed by Gemini (6.95), GPT-4 (5.80), and GPT-3.5 (3.71). This trend is similar when examining the performance playing against weak citizens. Note that these metrics yield different results regarding the best-performing model: WR indicates GPT-4 as the best, while LR identifies Llama.
This discrepancy is due to multiple winning conditions specified in the SDG. These conditions allow players to win games in several ways. Therefore, to accurately identify the skills that contribute to winning, it is necessary to consider multiple abilities simultaneously. However, since WR and LR consider only the final state of the game, these metrics cannot examine different abilities separately. For example, in SpyGame, LR indicates how well a spy survives during the game. However, a longer survival does not always guarantee victory, leading to potential discrepancies between LR and WR. Hence, a microscopic method is required for a comprehensive analysis of LLM’s performance in various facets.
TABLE II: Quantitative results for playing against strong citizens
(a) General (b) Subtext inference (c) Deceptive Control
Winning Living Guess Notice Information Information Caught Vote Vote
Model Rate (%) Round (#) Success (%) Rate (%) Catching Deduction (%) Rate (%) Rate (%) Entropy
Gemini 57.14 6.95 50.00 $3$ 30.00 022.22 35.7 0.348
(14) (20) (9) (42)
GPT-3.5 33.33 3.71 27.78 28.57 - 28.57 007.14 25.0 0.730
(18) (21) (14) (12)
GPT-4 80.95 5.80 81.25 61.90 - 61.90 025.00 33.3 0.953
(16) (21) (4) (30)
Llama2 52.38 9.00 - 09.52 - 00.00 100.00 45.2 0.634
(0) (19) (10) (126)
Total 55.95 6.365 35.78 $3$ 30.86 037.83 40.47 0.642
Metrics whose lower values indicate better results.
Note that Information Catching was not measured playing against strong citizens.
TABLE III: Quantitative results for playing against weak citizens
(a) General (b) Subtext inference (c) Deceptive control
Winning Living Guess Notice Information Information Caught Vote Vote
Model Rate (%) Round (#) Success (%) Rate (%) Catching (%) Deduction Rate (%) Rate (%) Entropy
Gemini 90.48 5.28 85.71 090.48 61.11 - 00.00 07.1 0.420
(14) (18) (2) (42)
GPT-3.5 71.43 2.52 71.43 076.19 92.86 - $0$
(21) (14) (6) (0)
GPT-4 90.48 2.42 90.48 100.00 87.50 - $0$
(21) (16) (2) (0)
Llama2 76.19 7.85 60.00 071.43 20.00 - 60.00 13.5 0.605
(5) (15) (5) (96)
Total 82.14 4.51 80.32 084.52 65.07 - 20.00 11.59 0.548
Metrics whose lower values indicate better results.
Note that Information Deduction was not measured playing against weak citizens.
IV-B Subtext Inference
We designed multiple metrics to enable a microscopic analysis of subtext inference. We operationally define subtext inference as the ability to uncover hidden information through deduction or by capturing pieces of information from the conversation. Subtext inference has two sub-abilities: information capture and information deduction. Information capture measures whether a spy notices and exploits critical hints inadvertently revealed by citizens. Since weak citizens generally leak such information unintentionally, information capture ability becomes essential for the spy to win the game against them. Since strong citizens rarely make such mistakes, it is highly challenging to observe information capture of a spy. On the other hand, information deduction mainly focuses on how effectively a spy infers a hidden location by combining various pieces of information. Since strong citizens generally conceal the key information, information deduction ability becomes essential for the spy to uncover information embodied in a citizen’s utterance.
To evaluate both information capture and deduction, we analyzed the outcomes of the spy’s guessing. The following subsections detail the design of the four metrics and their results. The results indicated that GPT-4 was the most powerful spy player, followed by Gemini, GPT-3.5, and Llama.
IV-B1 Metric design
To assess the spy’s ability to identify hidden intent, we propose four metrics based on their guessing behavior. These metrics range from relatively summative to more formative.
  • Guess Success (GS) measures how much the guessing of a spy contributed to their victory. This metric focuses only on games where the spy attempted to guess, rather than all played games. We calculate GS by dividing the number of games in which the spy made a successful guess by the number of games in which guessing was attempted. A higher GS indicates that the spy has successfully identified hidden locations, demonstrating strong subtext inference skills.
  • Notice Rate (NR) indicates how well a spy noticed the hidden location across all games. It is calculated by computing the percentage of games in which the spy correctly detected the hidden location over the total number of games played. Unlike GS, which is measured only when the spy decides to reveal their identity publicly, NR can be computed across all games. This is because NR mainly considers whether the spy notices the hidden location, regardless of whether they decide to guess.
  • Information Catching (IC) measures how often the spy successfully leverages exposed information to achieve a victory. Unlike NR, which only assesses the ability of the spy to identify the hidden location, IC evaluates the effectiveness in employing such information to win. We specifically focus on games where weak citizens directly exposed information about the hidden location instead of considering all the played games. Here, the exposure means a case where a citizen’s utterance literally contains the hidden location [^5] Qualitative analysis in Section V-B1 discusses the case.. We measured IC by computing the proportion of games in which the spy successfully guessed the hidden location from games where exposure occurred. Thus, a high IC score reflects the ability of the spy to achieve victory by effectively capturing and utilizing key information.
  • Information Deduction (ID) is similar to IC but it applies to games where exposure did not occur. This metric measures how well the spy combined pieces of information to determine the hidden location. To compute ID, we divided the number of games in which the spy successfully guessed the hidden location by the number of games without any exposure. Therefore, a high ID indicates that the spy successfully combined clues from the conversation and utilized them to achieve victory.
We provide a multifaceted interpretation of subtext inference by jointly analyzing IC and ID metrics. Based on whether each metric is above or below its average, we classify model behavior into three combinations. The first combination, when both IC and ID are high, represents an ideal scenario in which the spy successfully identifies the hidden location regardless of the type of citizen. The second combination occurs when only one of the metrics, either IC or ID, is higher than the average. This suggests a relative lack of proficiency in the corresponding sub-ability. For instance, a model with high IC but low ID demonstrates strong information capture but weak information deduction. The final and least favorable combination for subtext inference is when both IC and ID are low. This combination indicates that the spy failed to infer the hidden location in most games.
Moreover, by combining GS and NR, we can explore whether the spy effectively incorporates the incoming information to maintain their guess. The ideal combination is when both GS and NR are high. This indicates that the spy consistently maintained a correct guess after identifying the hidden location, until the end of the game. In contrast, a high GS paired with a low NR suggests that while the spy struggled to predict the hidden location initially, they managed to maintain their guess once the correct location was identified. Meanwhile, the combination of low GS and high NR implies that although the spy successfully guessed the hidden location, the spy is prone to discarding the guess easily. The worst-case scenario is when both NR and GS are low, indicating a significant failure in both identifying and maintaining the correct hidden location.
IV-B2 Result
Tables II (b) and III (b) present the measurements for subtext inference when playing against strong and weak citizens, respectively. The results indicate that GPT-4 has sufficient ability compared to the other models. GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and Llama (81.25, 50.0, 27.78, and 0%, respectively). The results in terms of NR show a similar trend: GPT-4 (61.90%), Gemini (33.33%), GPT-3.5 (28.57%), and Llama (9.52%). Finally, regarding ID, GPT-4 was the best model, followed by Gemini, GPT-3.5, and Llama (61.90, 30.0, 28.57, and 0%). In summary, GPT-4 demonstrated the best performance against strong citizens in subtext inference. Meanwhile, Llama showed the lowest performance against strong citizens.
Against weak citizens, GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and Llama (90.48, 85.71, 71.43, and 60%, respectively). NR displayed the same trend: GPT-4 (100%), Gemini (90.48%), GPT-3.5 (76.19%), and Llama (71.43%). Lastly, regarding IC, GPT-3.5 (92.86%) showed the best performance, differing in GS and NR, followed by GPT-4 (87.5%), Gemini (61.11%), and Llama (20.0%). Similar to the results against strong citizens, GPT-4 showcased good performance against weak citizens.
IV-B3 Discussion
In this section, we derive insights using the proposed metrics by conducting a multifaceted analysis. This discussion encompasses three aspects: (1) a complementary relationship between IC and ID; (2) an exploration of the spy’s filtering behavior on incoming information; and (3) the need for formative metrics concerning subtext inference.
First, IC and ID are complementary metrics that enable multifaceted analysis of subtext inference. As aforementioned, these two metrics can be combined to form three distinct partitions, each representing different levels of subtext inference ability. Using these partitions, we can determine the best model possessing sufficient sub-abilities for subtext inference; In our analysis, GPT-4 demonstrated high scores in both IC and ID, with its ID being double that of the second-place model. Meanwhile, GPT-3.5 falls within the high-IC region but not the high-ID region. This indicates that while GPT-3.5 struggles to combine hints from the conversation to deduce the hidden location.
Second, by analyzing GS and NR, we can examine how the spy filters incoming information to maintain their guess. As suggested in the metric design, we can use four quadrants made by GS and NR. Among the four models, GPT-4 belongs to the first partition and achieves a high score on both metrics. This indicates that GPT-4 can filter useful information, reinforcing the belief of the spy in the hidden location. This could be a hint that GPT-4 has strong information deduction sub-ability. Meanwhile, belonging to different partitions can induce differences in such filtering behaviors. For example, although Gemini and GPT-3.5 exhibited similar NR scores against strong citizens, GS of Gemini was almost double that of GPT-3.5. This difference suggests that Gemini is more selective in using guess actions than GPT-3.5, as evidenced by the fact that Gemini used 14 guess actions regardless of citizens, whereas GPT-3.5 utilized 21 (against weak) and 18 (against strong) guess actions.
Finally, our metrics regarding subtext inference enable more detailed analyses of the factors contributing to the scores of simple metrics. Let us consider the simple metric, Winning Rate (WR), on Gemini models. Against strong citizens, Gemini showed a similar WR to Llama. Meanwhile, against weak citizens, Gemini exhibited the same WR as GPT-4. However, these observations are somewhat shallow when we consider our metrics, as the three models display different subtext inference abilities. For example, against strong citizens, Gemini had a higher ID score than Llama in that Llama completely failed to combine clues about the hidden location. Similarly, against weak citizens, we can conclude that GPT-4 is slightly better than Gemini because of its higher GS and NR scores. Thus, proposed metrics can reveal finer differences between LLMs than existing metrics.
In addition, such a detailed analysis revealed that GPT-3.5 is overly confident in its deduction results. As illustrated in Tables II and III, the spy with GPT-3.5 demonstrated relatively lower NR scores; however, the spy used guessing actions more frequently. That is, the spy rushed into making guesses even when their inference about the hidden location was incorrect. One possible explanation is that GPT-3.5 may treat its own inference as hard evidence, as we will discuss in Section V. We suspect that this mechanism may have contributed to GPT-3.5 achieving the lowest WR score, despite its relatively good subtext inference ability.
IV-C Deceptive Control
In addition to subtext inference, we designed four metrics for the multifaceted analysis of deceptive control. Here, we operationally define deceptive control as the ability to blend in with others by controlling information. Deceptive control encompasses two sub-abilities: evading suspicion and dispersing suspicion. Evading suspicion mainly considers how well the spy avoids being suspected by others. It is quantified by the number of suspicions directed at the spy. On the other hand, dispersing suspicion primarily considers how difficult the spy makes it for all citizens to converge on a single suspect. It is quantified by the probability distribution of suspicions. To estimate the number and distribution of suspicions, we utilized logs of Day/Final votes. The following subsections describe the design of the four metrics and their results. The results indicate that GPT-3.5 is the most powerful spy player, followed by GPT-4, Gemini, and Llama.
IV-C1 Metric design
Based on day/final votes and accusations during a game, we suggest three metrics for measuring deceptive control, ranging from relatively summative to highly formative.
  • Caught Rate (CR) is how frequently the spy is caught by citizens in the Day/Final vote. We compute this by dividing the number of games in which the spy was caught by the number of games lost. Thus, a high CR indicates that the majority of citizens easily identified the spy player during the game. Note that this is a summative metric for the spy’s deceptive control ability because it only considers the final state of the game.
  • Vote Rate (VR) is the proportion of suspicions about the spy across all votes that occurred. We measure this by computing the percentage of votes the spy received from citizens across the total number of votes, excluding the spy’s votes. Therefore, a high VR indicates that the spy failed to evade suspicion during the game.
  • Vote Entropy (VE) is how scattered the suspicions about the spy are in the game. We measure this by computing the entropy of the votes that citizens casted for each game and averaging them. Note that in SpyGame, a player’s suspicion cannot affect another player’s suspicion because they cannot have a discussion together. Thus, when a spy successfully blends with citizens, citizens have difficulty narrowing down the number of spy candidates. Thus, the value of VE increases.
We devised a multifaceted interpretation of deceptive control by using VR and VE simultaneously. By partitioning VR and VE by their average values, we derived four combinations. First, the most balanced combination was a low VR and high VE. This is an ideal combination because the spy can successfully evade others’ suspicions and lead citizens to discourage each other from narrowing down spy candidates. Second, models in which VR and VE are both low or high lack one of the two sub-abilities. Finally, the worst combination occurs when VR is high and VE is low. This combination indicates that citizens’ suspicions were concentrated on the spy and that the spy failed to undermine suspicions.
IV-C2 Result
Tables II (c) and III (c) elaborate measurements of deceptive control when playing against strong and weak citizens, respectively. Note that some missing values for VR and VE indicate that the spy guessed the hidden location for all games, and no votes were conducted.
The results indicate that the variants of GPT have sufficient ability compared to other models. Against strong citizens, GPT-3.5 was the best model in terms of CR, followed by Gemini, GPT-4, and Llama (7.1, 22.2, 25.0, and 100%, respectively). The results for VR showed a similar tendency: GPT-3.5 (25%), GPT-4 (33.3%), Gemini (35.7%), and Llama (45.2%). Finally, regarding VE, GPT-4 was the best model, followed by GPT-3.5, Llama, and Gemini (0.953, 0.730, 0.634, and 0.348, respectively). Interestingly, GPT-3.5 showed deceptive control performance similar to that of GPT-4, although GPT-4 showed much higher performance in other language-based tasks[1].
Against weak citizens, GPT variants showed good performance compared to others. Because GPT variants ended the game by guessing, citizens had no chance to find the spy. Thus, CR values for GPT-3.5 and GPT-4 were zero, and VR and VE could not be measured using these models. Other models showed a tendency similar to that against strong citizens. Regarding CR and VR, Gemini (0% and 7.1%, respectively) was better than Llama (60% and 13.5%). However, regarding VE, Gemini (0.420) performed worse than Llama (0.605), as shown by results against strong citizens.
IV-C3 Discussion
In this section, we discuss how to derive insights using the proposed metrics to conduct a multifaceted analysis of deceptive control. This discussion includes the following two aspects: (1) a complementary relationship between Vote Rate (VR) and Vote Entropy (VE), and (2) additional perspectives captured by adopting formative metrics.
First, VR and VE are complementary metrics that enable a multifaceted analysis of deceptive control. As suggested above, these two metrics can form three partitions explaining the different levels of capacity regarding deceptive control. Using these partitions, we can determine the best model for balancing two sub-abilities of deceptive control; GPT family showed gameplay balanced between VR and VE. These models can evade and disperse suspicion to win the game. In contrast, the worst model in terms of deceptive control was Llama. Although Llama performed better against weak citizens than strong citizens, Llama player was often suspected as a spy and disregarded the fact that the majority of players doubted him/her. So, the two sub-abilities are both good or both bad in these two cases. Gemini showed a different behavior: the model was good at evading suspicion but struggled to disperse suspicion. Thus, as illustrated above, VR and VE can provide a wider picture of deceptive control behavior.
Second, proposed metrics provide additional perspectives on investigating deceptive control, compared to simple summative metrics. In section IV-A, we observed a discrepancy between two summative metrics, win rate (WR) and living round (LR). For example, both GPT-3.5 and Llama showed relatively low WR scores, but regarding LR, they achieved relatively low and high performance, respectively. So, using summative metrics only, we may conclude that Llama outperforms GPT-3.5 in SpyGame. However, the proposed metrics suggested different results; GPT-3.5 is good at controlling information, but Llama is not. This example shows that our proposed metrics are more suitable for examining deceptive control ability than existing summative metrics.
Similarly, our proposed metrics are helpful in identifying gameplay problems. For example, both GPT-4 and GPT-3.5 showed relatively low LR scores; however, WR achieved the best and worst scores, respectively. Therefore, using only summative metrics, we can conclude that GPT-4 outperforms GPT-3.5 in SpyGame. However, the proposed metrics suggest different results: both GPT-4 and GPT-3.5 are good at deceptive control regarding VR and VE.
V Qualitative Analysis
To reconfirm the findings of quantitative analysis from another perspective, we conducted a thematic analysis. We defined three types of reasoning errors in LLM reasoning while deciding their actions by coding abnormal patterns of 15 randomly selected gameplay logs. As a result, we found that these patterns are related to psychological concepts, including identity confusion, memory distortion, and dissociation[23]. In addition, we found that these concepts are related to the results of the quantitative analysis, especially for caught rate (CR) and guess success (GS). The following subsections illustrate the procedure and results of this qualitative analysis.
V-A Analysis Procedure
Through thematic analysis, we labeled the abnormal patterns of players in the gameplay logs. To enrich our quantitative analysis results, we adopted a bottom-up inductive approach to discover the underlying patterns in the data. The detailed analysis process was as follows. First, we defined the initial labels of the abnormal patterns from 15 randomly selected gameplay logs. Three annotators with computer science backgrounds participated in this process. They read the logs to understand the spy’s gameplay situation when establishing initial labels. To avoid bias towards a specific LLM, we hid the information of the models in the process. Second, we refined these initial labels through an iterative process. Because LLMs perform one or two reasoning steps to decide on an action during the gameplay of SpyGame, a gameplay log is a series of several reasoning steps. So, we used one step as the unit for this analysis. We repeated this process five times between December 29, 2023 and June 1, 2024 until agreement on the identified labels was saturated. The agreement was measured using Fleiss’ Kappa[24], and the final agreement between the annotators ranged from 0.486 to 0.646, which indicates moderate agreement. Finally, using these settled labels, we labeled 506 reasoning steps from 168 games. Consequently, we identified three major categories and five subcategories.
TABLE IV: Frequency of each category found in qualitative analysis
$N$* Character ambiguity Memory distortion Dissociation
Against weak citizens
Gemini 61 1.63 4.91 3.27
GPT-3.5 35 2.85 2.85 8.57
GPT-4 34 0.00 0.00 5.88
Llama2 77 9.09 6.49 9.90
Against strong citizens
Gemini 82 3.65 6.09 15.68
GPT-3.5 39 0.00 10.25 23.07
GPT-4 71 4.22 4.22 9.85
Llama2 107 9.09 10.28 14.95
* The total number of reasoning steps in each condition
V-B Resulting Categories
This section illustrates the results of our thematic analysis. For each category, we explain its definition, report its frequency, and provide some examples. Table IV summarizes the frequencies of all categories.
V-B1 Exposure
Exposure is a case in which a citizen’s utterance literally reveals the hidden location. The Fleiss’ Kappa value for this category was 0.7875. Note that this pattern can only be identified in utterances rather than in reasoning steps. Therefore, we report the frequency of this category based on utterances; among 729 citizen utterances, 125 were labeled as exposure. The following quote provides a representative example:
[The hidden location is an airplane]
Llama2 (Citizen, 1st-class passenger):
‘‘What is the maximum speed of an airplane in an emergency situation?’’
We suspect that LLMs may not have sufficient common sense to understand the game rules, causing this pattern. In particular, SpyGame has a rule “citizens should not expose the location to the spy.” A human citizen player can notice that a direct mention to other players of the hidden location during a game causes their defeat. However, some LLMs may not have this common sense among human players; they simply go through the mentions. Therefore, LLMs believe that they can reveal their information to other citizens because the game rule does not algorithmically prohibit such behaviors and the citizens already know that fact. Thus, citizen LLMs sometimes publicly reveal their hidden information.
V-B2 Dissociation
Dissociation is a case in which the spy employs their own utterances as evidence of the hidden location because they misunderstand the source of the utterance as other players. The annotators agreed on this category with a Fleiss’ Kappa value of 0.646. In this category, we identified 59 of 506 reasoning steps in the data. An example is as follows:
[Player4, the spy, tries to guess the hidden location.]
Llama2 (Spy): Based on the conversation, I would say that the location is most likely a workplace or office. Player4 described their environment as well-lit, well-ventilated, ...
We suspect that this category occurs because LLMs are not good at differentiating between numerical symbols. As previous studies have reported that LLMs suffer from biases induced by proper names[25], we named each player as ‘Player#’ (e.g., Player1) to avoid such biases. Therefore, when an LLM attempts to collect clues from the conversation log, it has to distinguish the source of each utterance using a single difference between player names: a single digit. However, previous studies have reported that LLMs struggle to recognize different digit tokens during their computation processes [26]. Thus, LLM players can misinterpret their words as those of others in the log, which triggers dissociation.
V-B3 Memory distortion
Memory distortion occurs when a spy uses false or distorted memory to decide their next action. For this category, the Fleiss’ Kappa value was 0.486 among the three annotators. We identified 32 steps with memory distortion among 506 reasoning steps from the data. An example is as follows:
[Previously, Player2 and Player3 did not have any conversation. The spy tries to guess the hidden location.]
GPT-3.5 (Spy): Based on the conversation between Player2 and Player3, it seems that the location is likely ...
We suspect this category is identical to the common “hallucination” phenomenon of LLMs, where a language model generates non-sensical or unfaithful text for the provided source input[27]. In a similar vein, our memory distortion makes the LLM use inappropriate, false, or distorted clues to decide the next action or utterance. These actions, which are affected by hallucination, can hinder the gameplay.
V-B4 Character ambiguity
Character ambiguity is a case in which the spy confuses their own identity with that of a citizen. This category is named after a similar psychological concept [28]. In an SDG, understanding a player’s identity is related to two major concepts affecting gameplay: (1) the team of the player and (2) the goal of the player. Therefore, we divided abnormal patterns of character ambiguity into two sub-categories: team misunderstanding and goal misunderstanding. These two subcategories can occur independently. We present details of the subcategories in the below paragraphs. In summary, Fleiss’ Kappa for this category was 0.563. Among 506 reasoning steps in the data, we identified 21 steps with character ambiguity.
Team misunderstanding
This is a case in which the spy disbelieves that they belong to citizens. Although this can occur in any reasoning step, we provide one example here:
[The spy tries to answer a citizen’s question. Before that, the spy generates its reasoning first.]
Gemini (Spy): Emphasize the importance of ambiance and personalized service, while avoiding giving away any specific details about the location...
Goal misunderstanding
This is a case in which the spy misbelieves that their goal is the same as that of the citizens. Here is an example:
[The spy tries to decide the next action. After selecting ‘question’ as the next action, the spy generates its reasoning.]
Llama2 (Spy): ... By asking this question, I hope to gain more information about the location and potentially identify the spy. If the spy is unaware of the restaurant’s cuisine, ...
We suspect that this category occurs is due to the memory limitation of LLMs. To understand a player’s identity correctly, an LLM should retain information about the player’s identity during the decoding process. Since we provided the game rule, identity of a player, and conversation logs together as an input prompt, the LLM needs to distinguish information about the identity from the input. However, such a distinction is difficult for the LLM because the other parts of the input prompt are much larger than the identity part. For example, the explanation of the game rules takes more than 130 words, whereas that of identity takes only 10 words. Thus, LLM players can confuse their identities with other characters written in the game rules, triggering character ambiguity.
V-C Reconfirmation of Quantitative Analysis
Combining the results of the two analyses confirms the findings of the quantitative analysis. Specifically, we discovered that these quantitative labels, except exposure, can be connected to two abilities in the quantitative analysis: subtext inference and deceptive control.
First, memory distortion and dissociation may hinder a spy from inferring subtext during the game. When we compared the frequency of memory distortion and dissociation with the results of the four subtext inference metrics, the trend of frequency was similar to that of the metrics. For example, against strong citizens, GPT-3.5 showed the highest errors in both memory distortion and dissociation and achieved the lowest performance in subtext inference. Similarly, GPT-4 showed fewer errors while achieving a higher performance than the other models. We suspect that this similarity between trends is due to the quality of information that the spy considers during their reasoning process. The spy player should select and combine pieces of information from the conversation to identify the hidden location. When memory distortion occurs, the spy can regard misinterpreted pieces as appropriate. Similarly, when dissociation occurs, the spy regards apocryphal information that they previously said as information provided by citizens. Therefore, such misunderstood pieces can be obstacles to reaching the correct answer.
Second, character ambiguity increases the likelihood of a spy’s failure of deceptive control. When we compared the frequency of character ambiguity and the results of the three metrics regarding deceptive control, the trend of frequency was similar to that of the metrics. For example, against strong citizens, GPT-3.5 showed the lowest errors in character ambiguity and achieved the highest performance in such metrics. Also, Llama showed more errors while achieving a lower performance than the other models. We suspect that this similarity between the trends is due to the loss of the spy’s need for deception. When character ambiguity occurs in the spy’s reasoning steps, the spy confuses their character with that of a citizen. Such confusion leads the spy to make an unusual utterance, which draws suspicion.
VI Conclusion
To assess whether LLMs can effectively support obscured communication, we addressed two key limitations in prior evaluation methodologies. The first is the reliance on coarse-grained quantitative metrics, which fail to capture fine-grained communicative behaviors. The second is the lack of structured qualitative analyses that can contextualize evaluation results or uncover the reasoning processes underlying model behavior. To investigate these limitations in a controlled setting, we conducted experiments using SpyGame, a variant of Spyfall, and collected gameplay logs from four LLMs. We analyzed these logs through both quantitative and qualitative methods. First, we proposed eight fine-grained metrics designed to evaluate two core abilities: subtext inference and deceptive control. These metrics demonstrated more effective in identifying subtle behavioral differences between models. Second, we performed a thematic analysis to identify reasoning failures not captured by quantitative metrics. This analysis revealed four major categories of failure that align with the patterns observed in the quantitative evaluation, offering complementary insights. Our findings highlight the importance of microscopic, multi-perspective evaluation for understanding how LLMs perform in obscured communicative settings. We hope this study contributes to a deeper understanding on LLMs in complex social interactions.